y=(sinx+cosx/sinx-cosx)then find dy/dx at x=15

Grade:Upto college level

1 Answers

Vinti Singhai
10 years ago
Given expression can be written as -(1+tanx)/(1-tanx)=-tan(x+pi/4)Hence dy/dx = -sec^2(x+pi/4)Vinti SinghaiaskIITians Faculty,B.E(Honors) in Elec &Comm, Gold Medalist